# Optimizing Layout of Recursive Datatypes with Marmoset

### v2

@@ 5.4.4 Globally optimizing multiple functions
Table 4: Runtime in seconds for different blog traversals when we use Marmoset to globally optimize the data layout. The input parameters are different from the single function case.
(a) Filter Blogs
Gibbon Mgreedy Msolver
Mean 1.926 918 888 888 888 81.92691888888888881.926\,918\,888\,888\,888\,81.926 918 888 888 888 8 0.101 968 213 131 313 130.101968213131313130.101\,968\,213\,131\,313\,130.101 968 213 131 313 13 0.104 246 677 777 777 770.104246677777777770.104\,246\,677\,777\,777\,770.104 246 677 777 777 77
Median 1.925 581.925581.925\,581.925 58 0.101 113 40.10111340.101\,113\,40.101 113 4 0.102 765 90.10276590.102\,765\,90.102 765 9
ub 1.930 799 059 311 418 21.93079905931141821.930\,799\,059\,311\,418\,21.930 799 059 311 418 2 0.102 606 986 857 571 250.102606986857571250.102\,606\,986\,857\,571\,250.102 606 986 857 571 25 0.105 125 781 081 590 640.105125781081590640.105\,125\,781\,081\,590\,640.105 125 781 081 590 64
lb 1.923 038 718 466 359 51.92303871846635951.923\,038\,718\,466\,359\,51.923 038 718 466 359 5 0.101 329 439 405 055 010.101329439405055010.101\,329\,439\,405\,055\,010.101 329 439 405 055 01 0.103 367 574 473 964 90.10336757447396490.103\,367\,574\,473\,964\,90.103 367 574 473 964 9
(b) Emph Content
Gibbon Mgreedy Msolver
Mean 1.582 479 656 565 656 51.58247965656565651.582\,479\,656\,565\,656\,51.582 479 656 565 656 5 1.388 232 272 727 272 71.38823227272727271.388\,232\,272\,727\,272\,71.388 232 272 727 272 7 1.325 701 141 414 141 51.32570114141414151.325\,701\,141\,414\,141\,51.325 701 141 414 141 5
Median 1.578 6781.5786781.578\,6781.578 678 1.384 181.384181.384\,181.384 18 1.320 4681.3204681.320\,4681.320 468
ub 1.589 285 432 654 959 71.58928543265495971.589\,285\,432\,654\,959\,71.589 285 432 654 959 7 1.396 201 464 419 347 41.39620146441934741.396\,201\,464\,419\,347\,41.396 201 464 419 347 4 1.336 001 926 809 490 71.33600192680949071.336\,001\,926\,809\,490\,71.336 001 926 809 490 7
lb 1.575 673 880 476 353 31.57567388047635331.575\,673\,880\,476\,353\,31.575 673 880 476 353 3 1.380 263 081 035 1981.3802630810351981.380\,263\,081\,035\,1981.380 263 081 035 198 1.315 400 356 018 792 31.31540035601879231.315\,400\,356\,018\,792\,31.315 400 356 018 792 3
(c) Tag Search
Gibbon Mgreedy Msolver
Mean 2.186 549 565 656 565 42.18654956565656542.186\,549\,565\,656\,565\,42.186 549 565 656 565 4 1.830 526 727 272 727 51.83052672727272751.830\,526\,727\,272\,727\,51.830 526 727 272 727 5 2.356 293 929 292 929 22.35629392929292922.356\,293\,929\,292\,929\,22.356 293 929 292 929 2
Median 2.182 5412.1825412.182\,5412.182 541 1.826 9591.8269591.826\,9591.826 959 2.351 9892.3519892.351\,9892.351 989
ub 2.193 582 984 981 6582.1935829849816582.193\,582\,984\,981\,6582.193 582 984 981 658 1.838 355 525 519 980 81.83835552551998081.838\,355\,525\,519\,980\,81.838 355 525 519 980 8 2.364 621 853 531 670 82.36462185353167082.364\,621\,853\,531\,670\,82.364 621 853 531 670 8
lb 2.179 516 146 331 472 72.17951614633147272.179\,516\,146\,331\,472\,72.179 516 146 331 472 7 1.822 697 929 025 474 11.82269792902547411.822\,697\,929\,025\,474\,11.822 697 929 025 474 1 2.347 966 005 054 187 72.34796600505418772.347\,966\,005\,054\,187\,72.347 966 005 054 187 7
@@ 5.4.5 Comparison of Marmoset against MLton
Figure 9: Performance comparison of Msolver with MLton. The tuple on top of each bar shows the median runtime in seconds of Msolver followed by the corresponding layout when compiled with MLton.
@@ 5.5 Cache behavior
Gibbon Marmoset
@@ 5.6 Tradeoffs between Marmoset’s solver and greedy optimization
(a) Filter blogs traversal
(b) Content Search traversal
(c) Tag Search traversal
Figure 10: Average compile times (99 runs) in (s) for different layouts and traversal combinations when compiled with Gibbon and when optimized by Mgreedy and Msolver. The bar in Blue is the time for Msolver after subtracting the time for solver IO and the bar in Red is the time taken to do the IO call to the DOCPLEX solver. The total compile time for Msolver includes the time for doing the CFG analysis, generating the access graph, the solver time, reordering the datatype and repairing the code, The majority of which is added by the solver IO.
We show the compile times in Figure 10 for different layout and traversal combinations when compiled with Gibbon, Mgreedy and Msolver respectively as a measure of relative costs. The compile times for Msolver include the time to generate the control flow graph, the field access graph, the solver time and the time to re-order the datatype in the code. The solver times are in the order of the number of fields in a data constructor and not the program size. Hence, the solver adds relatively low overhead. Since the compiler does an IO call to the python solver, there is room for improvement in the future to lower these times. For instance, we could directly perform FFI calls to the CPLEX solver by lowering the constraints to C code. This would be faster and safer than the current implementation. In addition, during the global optimization, we call the solver on each data constructor as of the moment, we could further optimize this by sending constraints for all data constructors at once and doing just one solver call.
Although the cost of Marmoset’s solver based optimization is higher that the greedy approach, it is a complementary approach which may help the user find a better layout at the cost of compile time. On the other hand, if the user wishes to optimize for the compile time, they should use the greedy heuristic.
@@ 5.7 Discussion: Scale of Evaluation
6 Discussion
Marmoset allows the user to provide optional manual constraints on the layout as either relative or absolute constraints through ANN annotation pragmas. The relative constraint allows the user to specify if a field A comes immediately after field B. The absolute constraint specifies an exact index in the layout for a field. Such pragmas may be beneficial if the user requires a specific configuration over data types or has information about performance bottlenecks.
@@ 6 Discussion
We could also look at a scenario where we optimize each function locally and use “shim” functions that copy one layout to another (the one required by the next function in the pipeline). Although the cost of copying may be high, it warrants further investigation. Areas of improvement purely on the implementation side include optimizing whether Marmoset dereferences a pointer to get to a field or uses the end-witness information as mentioned in section 2. Lesser pointer dereferencing can lower instruction counts and impact performance positively. We would also like to optimize the solver times as mentioned in 10.
@@ 7.1 Cache-conscious data
  • ccmorph: This works on tree-like data structures, and it relies on the programmer making a calculated guess about the safety of the operation on the tree-like data structure. It performs two major optimizations: clustering and coloring. Clustering take a the tree like data structure and attempts to pack likely to be accessed elements in the structure within the same cache block. There are various ways to pack a subtree, including clustering $k$ nodes in a subtree together, depth first clustering, etc. Coloring attempts to map simultaneously accessed data elements to non-conflicting regions of the cache.
  • ccmalloc: This is a memory allocator similar to malloc which takes an additional parameter that points to an existing data structure element which is likely accessed simultaneously. This requires programmer knowledge and effort in recognizing and then modifying the code with such a data element. ccmalloc tries to allocate the new data element as close to the existing data as possible, with the initial attempt being to allocate in the same cache block. It tries to put likely accessed elements on the same page in an attempt to improve TLB performance.
Franco et al. [13, 14] suggest that the layout of a data structure should be defined once at the point of initialization, and all further code that interacts with the structure should be “layout agnostic”. Ideally, this means that performance improvements involving layout changes can be made without requiring changes to program logic. To achieve this, classes are extended to support different laouts, and types carry layout information—code that operates on objects may be polymorphic over the layout details.
@@ 8 Conclusions
This paper introduced Marmoset, which builds on Gibbon to generate efficient orders for Algebraic Datatypes. We observed that a fairly straightforward control-flow and data-flow analysis allows Marmoset to identify opportunities to place fields of a data constructor near each other in memory to promote efficient consecutive access to those fields. Because a given function might use many fields in many different ways, Marmoset adopts an approach of formulating the data layout problem as an ILP, with a cost model that assigns an abstract cost to a chosen layout. This ILP can be further constrained with user-provided layout guidelines. Armed with problem formulation, an off-the-shelf ILP solver allows Marmoset to generate minimal-(abstract)-cost layouts for algebraic datatypes. Marmoset then uses these layouts to synthesize a new ADT and prior work’s Gibbon compiler toolchain to lower the code into an efficient program that operates over packed datatypes with minimal pointer chasing.
We found, across a number of benchmarks, that Marmoset is able to effectively and consistently find the optimal data layout for a given combination of traversal function and ADT. We showed, experimentally, that Marmoset-optimized layouts significantly outperform not only Gibbon’s default layouts, but also GHC’s standard compiler.

### v3

@@ 5.4.1 Globally optimizing multiple functions
Table 5: Run-time mean and 95% confidence interval (ub, lb) for the blog software benchmarks when Marmoset optimizes the data layout globally (seconds). The input parameters are different from the single-function optimization case.
Benchmark name Gibbon Marmoset
hiadctb Mgreedy Msolver
FilterBlogs 2.230 235 366 197 1832.2302353661971832.230\,235\,366\,197\,1832.230 235 366 197 183 (2.229 660 184 321 464 32.22966018432146432.229\,660\,184\,321\,464\,32.229 660 184 321 464 3, 2.230 810 548 072 9022.2308105480729022.230\,810\,548\,072\,9022.230 810 548 072 902) 0.111 555 726 760 563 380.111555726760563380.111\,555\,726\,760\,563\,380.111 555 726 760 563 38 (0.111 297 542 687 853 760.111297542687853760.111\,297\,542\,687\,853\,760.111 297 542 687 853 76, 0.111 813 910 833 273 010.111813910833273010.111\,813\,910\,833\,273\,010.111 813 910 833 273 01) 0.095 000 133 239 436 610.095000133239436610.095\,000\,133\,239\,436\,610.095 000 133 239 436 61 (0.094 963 564 634 950 20.09496356463495020.094\,963\,564\,634\,950\,20.094 963 564 634 950 2, 0.095 036 701 843 923 020.095036701843923020.095\,036\,701\,843\,923\,020.095 036 701 843 923 02)
EmphContent 1.574 749 915 492 957 41.57474991549295741.574\,749\,915\,492\,957\,41.574 749 915 492 957 4 (1.574 266 853 063 465 61.57426685306346561.574\,266\,853\,063\,465\,61.574 266 853 063 465 6, 1.575 232 977 922 449 21.57523297792244921.575\,232\,977\,922\,449\,21.575 232 977 922 449 2) 1.383 689 154 929 577 31.38368915492957731.383\,689\,154\,929\,577\,31.383 689 154 929 577 3 (1.383 560 188 249 942 41.38356018824994241.383\,560\,188\,249\,942\,41.383 560 188 249 942 4, 1.383 818 121 609 212 21.38381812160921221.383\,818\,121\,609\,212\,21.383 818 121 609 212 2) 1.322 207 957 746 478 71.32220795774647871.322\,207\,957\,746\,478\,71.322 207 957 746 478 7 (1.322 079 854 772 291 41.32207985477229141.322\,079\,854\,772\,291\,41.322 079 854 772 291 4, 1.322 336 060 720 6661.3223360607206661.322\,336\,060\,720\,6661.322 336 060 720 666)
TagSearch 2.196 739 676 056 3392.1967396760563392.196\,739\,676\,056\,3392.196 739 676 056 339 (2.196 016 148 977 5722.1960161489775722.196\,016\,148\,977\,5722.196 016 148 977 572, 2.197 463 203 135 1062.1974632031351062.197\,463\,203\,135\,1062.197 463 203 135 106) 1.830 104 352 112 6761.8301043521126761.830\,104\,352\,112\,6761.830 104 352 112 676 (1.828 666 212 218 7451.8286662122187451.828\,666\,212\,218\,7451.828 666 212 218 745, 1.831 542 492 006 607 21.83154249200660721.831\,542\,492\,006\,607\,21.831 542 492 006 607 2) 2.350 361 915 492 957 42.35036191549295742.350\,361\,915\,492\,957\,42.350 361 915 492 957 4 (2.350 152 096 698 4252.3501520966984252.350\,152\,096\,698\,4252.350 152 096 698 425, 2.350 571 734 287 492.350571734287492.350\,571\,734\,287\,492.350 571 734 287 49)
FilterBlogs
EmphContent
TagSearch
Figure 9: Performance comparison of Msolver with MLton. The pair of numbers on top of each bar shows the median runtime in seconds of Msolver followed by the corresponding layout when compiled with MLton.
FilterBlogs
EmphContent
TagSearch
Figure 10: Average compile times (99 runs) in seconds for different layouts and traversal combinations when compiled with Gibbon and when optimized by Mgreedy and Msolver.
@@ 5.4.2 Comparison of Marmoset against MLton
5.4.3 Comparison of Marmoset against Ghc
We compare Marmoset’s performance to GHC, which is especially optimized to run functional programs.
@@ 5.4.3 Comparison of Marmoset against Ghc
Figure 11: Performance comparison of Msolver with Ghc.
Figure 11 shows the the speedup of Marmoset over Ghc when evaluated strictly for an apples to apples comparison. We pass the optimization option -O2 to Ghc while compiling. As shown, the performance of Marmoset is better than Ghc by significant margins for all the layouts and traversals. Since the types in Ghc are boxed, such a behavior is expected because it adds more instructions (pointer de-referencing) and results in worse spatial locality.
@@ 5.5 Cache behavior
Benchmark name or metric Gibbon Marmoset
@@ 5.6 Tradeoffs between Marmoset’s solver and greedy optimization
In Figure 10 (p. 10), we show the compile times for different layout and traversal combinations when compiled with Gibbon, Mgreedy and Msolver respectively as a measure of relative costs. The compile times for Msolver include the time to generate the control flow graph, the field access graph, the solver time and the time to re-order the datatype in the code. The solver times are in the order of the number of fields in a data constructor and not the program size. Hence, the solver adds relatively low overhead. Since the compiler does an IO call to the python solver, there is room for improvement in the future to lower these times. For instance, we could directly perform FFI calls to the CPLEX solver by lowering the constraints to C code. This would be faster and safer than the current implementation. In addition, during the global optimization, we call the solver on each data constructor as of the moment, we could further optimize this by sending constraints for all data constructors at once and doing just one solver call.
Although the cost of Marmoset’s solver based optimization is higher than the greedy approach, it is a complementary approach which may help the user find a better layout at the cost of compile time. On the other hand, if the user wishes to optimize for the compile time, they should use the greedy heuristic.
@@ 5.7 Discussion: Scale of Evaluation
6 Future Work
Marmoset could allow the user to provide optional constraints on the layout (either relative or absolute) through pragmas. A relative constraint would allow the user to specify if a field A comes immediately after field B. An absolute constraint would specify an exact index in the layout for a field. Such pragmas may be useful if the user requires a specific configuration of a data type for external reasons or has information about performance bottlenecks.
@@ 6 Future Work
We could also look at a scenario where we optimize each function locally and use “shim” functions that copy one layout to another (the one required by the next function in the pipeline). Although the cost of copying may be high, it warrants further investigation. Areas of improvement purely on the implementation side include optimizing whether Marmoset dereferences a pointer to get to a field or uses the end-witness information as mentioned in section 2. Lesser pointer dereferencing can lower instruction counts and impact performance positively. We would also like to optimize the solver times as mentioned in 5.6.
@@ 7.1 Cache-conscious data
ccmorph
This works on tree-like data structures, and it relies on the programmer making a calculated guess about the safety of the operation on the tree-like data structure. It performs two major optimizations: clustering and coloring. Clustering take a the tree like data structure and attempts to pack likely to be accessed elements in the structure within the same cache block. There are various ways to pack a subtree, including clustering $k$ nodes in a subtree together, depth first clustering, etc. Coloring attempts to map simultaneously accessed data elements to non-conflicting regions of the cache.
ccmalloc
This is a memory allocator similar to malloc which takes an additional parameter that points to an existing data structure element which is likely accessed simultaneously. This requires programmer knowledge and effort in recognizing and then modifying the code with such a data element. ccmalloc tries to allocate the new data element as close to the existing data as possible, with the initial attempt being to allocate in the same cache block. It tries to put likely accessed elements on the same page in an attempt to improve TLB performance.
Franco et al. [13, 14] suggest that the layout of a data structure should be defined once at the point of initialization, and all further code that interacts with the structure should be “layout agnostic”. Ideally, this means that performance improvements involving layout changes can be made without requiring changes to program logic. To achieve this, classes are extended to support different layouts, and types carry layout information—code that operates on objects may be polymorphic over the layout details.
@@ 8 Conclusions
This paper introduces Marmoset, which builds on Gibbon to generate efficient orders for algebraic datatypes. We show that a straightforward control-flow and data-flow analysis allows Marmoset to identify opportunities to place fields of a data constructor near each other in memory to promote efficient consecutive access to those fields. Because a given function might use many fields in many different ways, Marmoset adopts an approach of formulating the data layout problem as an ILP, with a cost model that assigns an abstract cost to a chosen layout. Armed with the ILP problem formulation, an off-the-shelf ILP solver allows Marmoset to generate minimal-(abstract)-cost layout for algebraic datatypes. Marmoset then uses the best layout to synthesize a new ADT and the Gibbon compiler toolchain to lower the code into an efficient program that operates over packed datatypes with minimal pointer chasing.
We show, across a number of benchmarks, that Marmoset is able to effectively and consistently find the optimal data layout for a given combination of traversal function and ADT. In our experiments, Marmoset-optimized layouts outperform not only Gibbon’s default layouts but also the popular SML compiler MLton.